OpenStudy (anonymous):

Explain why the quantum number set (3, 3, -2, +½) is not possible for an electron in a ground-state atom.

OpenStudy (anonymous):

um did you google it ?

OpenStudy (anonymous):

yea i googled it. I just don't really know the answer.

OpenStudy (anonymous):

do you know the answer ? @sarah786

OpenStudy (anonymous):

I don't actually .. I read it but a long time ago .. did yahoo answer not statisfied yo ?

OpenStudy (anonymous):

When i typed the question in google, there was nothing that came up. @sarah786

OpenStudy (anonymous):

Its easiy But in typing ... such a waste of time .. its exactly what that given on yahoo.answer .. I post web to you .I hope this helps U :)

OpenStudy (anonymous):

oh ok thank you! @sarah786

OpenStudy (anonymous):

Mention not ..thanx for medal ..Actually in real sense I couldn't help u :( But anyways thanxx for medal :)

OpenStudy (wolfe8):

OK I'm thinking, maybe because the first(n) and second (l) are the same. L should be n-1. That's what I think

OpenStudy (anonymous):

no i figured it out: The quantum numbers tell you where an electron is in ground-state atom. The format is (n, L, mL, ms) so your n=3 (ok) L=2 (ok). if n=3, L can= 0, 1 or 2 mL=-3 (wrong) if L=2, mL can only be -2 -1 0 +1 or +2 ms=+1/2 (ok) @wolfe8

OpenStudy (wolfe8):

Well yea basically what I said too. l cannot be the same as n. And btw, that link is for a question that has a different set of quantum numbers. Just noting.